mymommy:

the number of cups of strawberries and a shortcake recipe is proportional to the number of cubs sugar in a recipe the recipe uses five cups of strawberry / 1 and 1/2 cup of sugar what is the context of portion ability for the relationship of cubs for strawberries to cups of sugar

Ghostmask:

number of cups of strawberries in a shortcake recipe is proportional to the number of cups of sugar in the recipe Let number of cups of strawberries = x number of cups of sugar = y x = ky Where k = constant of proportionality number of cups of strawberries = 5 number of cups of sugar = 1/4 x = ky 5 = k * 1/4 5 = 1/4k Divide both sides by 1/4 k = 5 ÷ 1/4 = 5 × 4/1 = 20 k = 20 Constant of proportionality, k = 20

mymommy:

ok thanks

Ghostmask:

np